OpenStudy (anonymous):

. Find the equation of a line which passes through point (-4, 5) and is parallel to the line formed by the equation y = 3x + 1. Express it in slope-intercept form.

OpenStudy (anonymous):

since the line is parallel to y= 3x+1, it will have the same slope, so m=3 \[y-y_{0}=m(x-x_{0}), y_{0}= 5, x_{0}=-4\] \[y-5=3(x-(-4))\rightarrow y=3(x-(-4))+5 \rightarrow y=3x+17\]

OpenStudy (curry):

answer ed it already

OpenStudy (curry):

in earlier posting

OpenStudy (curry):

nadeem help me on on me